OpenStudy (anonymous):

(x/sqrt(x^2+1)+1)/(x+sqrt(x^2+1)) simplify (x/sqrt(x^2+1)+1)/(x+sqrt(x^2+1)) simplify @Mathematics

OpenStudy (anonymous):

plz show steps

OpenStudy (anonymous):

type using equation tool

myininaya (myininaya):

\[\frac{\frac{x}{\sqrt{x^2+1}+1}}{x+\sqrt{x^2+1}}\]?

OpenStudy (anonymous):

yeah it looks like that

OpenStudy (anonymous):

except the 1 is outside the squareroot

OpenStudy (anonymous):

numerator

myininaya (myininaya):

so both ones?

OpenStudy (anonymous):

Is it sqrt for both the ones?

myininaya (myininaya):

here you type go to equation button below and for fraction do frac{something}{stuff} and if you want square root of something do sqrt{something}

myininaya (myininaya):

\[\frac{\frac{x}{\sqrt{x^2+1}}+1}{x+\sqrt{x^2+1}}\]

myininaya (myininaya):

multiply top and bottom by sqrt(x^2+1)

OpenStudy (anonymous):

\[(x+1)^{-1}\]

myininaya (myininaya):

\[\frac{\frac{x}{\sqrt{x^2+1}}+1}{x+\sqrt{x^2+1}} \cdot \frac{\sqrt{x^2+1}}{\sqrt{x^2+1}}\]

OpenStudy (anonymous):

ok

myininaya (myininaya):

\[\frac{x+\sqrt{x^2+1}}{x \sqrt{x^2+1}+(x^2+1)}\]

OpenStudy (anonymous):

im following

myininaya (myininaya):

\[\frac{x+\sqrt{x^2+1}}{x \sqrt{x^2+1}+(x^2+1)} \cdot \frac{x \sqrt{x^2+1}-(x^2+1)}{x \sqrt{x^2+1}-(x^2+1)}\] \[\frac{(x+\sqrt{x^2+1})(x \sqrt{x^2+1}-(x^2+1))}{x^2 (x^2+1)-(x^2+1)^2}\] \[=\frac{(x+\sqrt{x^2+1})(x \sqrt{x^2+1}-(x^2+1))}{x^4+x^2-(x^4+2x^2+1)}\] \[=\frac{(x+\sqrt{x^2+1})(x \sqrt{x^2+1}-(x^2+1))}{-x^2-1}\] \[=\frac{(x+\sqrt{x^2+1})(x \sqrt{x^2+1}-(x^2+1))}{-(x^2+1)}\]

myininaya (myininaya):

\[=\frac{x^2 \sqrt{x^2+1}-x(x^2+1)+x(x^2+1)-(x^2+1)^{\frac{1}{2}+1}}{-(x^2+1)}\] \[=\frac{x^2 \sqrt{x^2+1}-(x^2+1)^\frac{3}{2}}{-(x^2+1)}\]

myininaya (myininaya):

\[=\frac{\sqrt{x^2+1}(x^2-(x^2+1))}{-(x^2+1)}=\frac{\sqrt{x^2+1}(-1)}{-(x^2+1)}\]

OpenStudy (anonymous):

ok it makes sense

myininaya (myininaya):

\[=\frac{\sqrt{x^2+1}}{x^2+1}\]

myininaya (myininaya):

does it really i did a lot of stuff you followed that? if so thats real good

OpenStudy (anonymous):

im double checking

myininaya (myininaya):

ok i did multiply top and bottom by the conjugate of the bottom to rationalize the bottom did you notice that?

myininaya (myininaya):

you know after i got rid of the compound fraction action

OpenStudy (anonymous):

yeah i got those parts

OpenStudy (anonymous):

i just wanted to compare with my answer

myininaya (myininaya):

ok

myininaya (myininaya):

i actually had to think on can i simplify this more after i got to a certain point i like paused for 5 seconds lol

OpenStudy (anonymous):

i know what you mean

OpenStudy (anonymous):

my paper right now is filled with a bunch of unnecesary moves

OpenStudy (anonymous):

i made it worse than the original statement

myininaya (myininaya):

sometimes it takes unnecessary steps to see you are going to a deadend

OpenStudy (anonymous):

yeah i know what you mean, but i really messed up this problem but i got a different answer.

myininaya (myininaya):

well i feel confident in my answer

OpenStudy (anonymous):

whops wrong equation

myininaya (myininaya):

hey i'm going to go to bed

myininaya (myininaya):

just make a new thread k?

OpenStudy (anonymous):

thx alot 4 your help

myininaya (myininaya):

np

OpenStudy (anonymous):

btw how do u get achievements?

myininaya (myininaya):

thats a good question i don't see the page that tells you what you need to do in order to receive achievement awards

myininaya (myininaya):

post it in the feedback section one of the admins will tell you

myininaya (myininaya):

you might have to wait a day for a reply

OpenStudy (anonymous):

ive been here for three days and answered 100s of questions

OpenStudy (anonymous):

and went up to lvl 13, its kinda like a game lol

OpenStudy (anonymous):

i enjoy helping though

myininaya (myininaya):

an addictive game

OpenStudy (anonymous):

a challenge at times

myininaya (myininaya):

ok off to brush my teeth

OpenStudy (anonymous):

the easy questions aren't really worth answering though,